OpenStudy (anonymous):

starting now, jordan will invest $250 into an ordinary annuity at the end of each monthly period, for a total of 30 years. assume that the annuity has an interest rate of 4.9 a) How much will he have invested by the end of the 30 years? b) how much interest will he have earned after the 30 years?

OpenStudy (wolf1728):

The amount invested is fairly easy 250 / month and there's 12 months per year for 30 years = $90,000.00

OpenStudy (wolf1728):

The annuity formula is attached. Total = monthly Amt * ([(1+r)^(n+1)-1] / r) -monthly amount n = months r is monthly rate Total = 250*([(1.0040833333)^361 -1] / 0.0040833333) -250 Total = 250*(3.3539431349 / .0040833333) -250 Total = 250*(821.3738356602) -250 Total = 205,343.46 -250 Total = 205,093.46 Interest Earned = 205,093.46 -90,000.00 Interest Earned = 115,343.46